# Power Battery Detection

### v1

@@ V-B Comparison with Corner Detection, Crowd Counting, General and Tiny Object Detection Solutions
Figure 13: Visual comparison with other general/tiny object detection-based [41, 40, 36, 38, 35], counting-based [53, 43, 49], corner detection-based [31, 30, 29] solutions. We directly visualize the predicted results (Ours: Segmentation map, General/Tiny object detection methods: Bounding box, Counting methods: Density map, Corner detection methods: Corner map) without any post-processing operations. Black dashed boxes ([--]) highlight the regions with obvious prediction failures. Best viewed zoomed in.
TABLE IV: Quantitative comparison of image segmentation methods. All ten metrics are widely used across various segmentation branches [67, 65, 98, 80], with details introduced in the survey [61]. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively. Best scores are highlighted in red. MDCNeXt’s relative improvements over MDCNet are highlighted in green. “Average” denotes the mean scores across Regular, Difficult, and Tough test splits. For prompt-based generalist models (i.e., SAM 2 and SegGPT), we use the same pure battery images and their corresponding masks as those in MDCNeXt as prompts to generate predictions for the test set.
Regular SAM 2 [78] ICLR 2025 Hiera-B+ [100] 0.9802 0.2473 0.2213 0.1979 0.5932 0.7026 0.5121 0.2722 0.3541 0.0177
SegGPT [76] ICCV 2023 ViT-L [19] 0.9818 0.2163 0.1958 0.1932 0.5541 0.7319 0.5551 0.1958 0.3880 0.0182
DeepLabV3+ [54] ECCV 2018 ResNet-101 [95] 0.9867 0.5078 0.4278 0.4106 0.6438 0.9124 0.5920 0.4379 0.2711 0.0147
SegFormer [67] NeurIPS 2021 MiT-B5 [67] 0.9848 0.5435 0.4326 0.4020 0.6726 0.9074 0.6020 0.4541 0.2820 0.0134
Spider [80] ICML 2024 ConvNeXt-L [101] 0.9888 0.6425 0.5970 0.6212 0.7836 0.9425 0.6820 0.6278 0.1422 0.0097
ZoomNeXt [102] TPAMI 2024 ResNet-50 [95] 0.9888 0.6320 0.6073 0.6320 0.7825 0.9617 0.6923 0.6342 0.1302 0.0092
MDCNet [21] CVPR 2024 ResNet-50 [95] 0.9909 0.6730 0.6635 0.7083 0.7912 0.9687 0.7432 0.7023 0.1107 0.0081
MDCNeXt ResNet-50 [95] 0.9924 0.6979 0.6971 0.7337 0.8113 0.9786 0.7857 0.7248 0.1101 0.0076
Difficult SAM 2 [78] ICLR 2025 Hiera-B+ [100] 0.9836 0.1989 0.1782 0.1678 0.5029 0.7302 0.5780 0.1011 0.4474 0.0187
SegGPT [76] ICCV 2023 ViT-L [19] 0.9817 0.1499 0.1359 0.1337 0.5281 0.6867 0.5335 0.1347 0.4233 0.0182
DeepLabV3+ [54] ECCV 2018 ResNet-101 [95] 0.9865 0.3691 0.3742 0.3542 0.4109 0.9102 0.5532 0.3801 0.3080 0.0142
SegFormer [67] NeurIPS 2021 MiT-B5 [67] 0.9888 0.3390 0.3512 0.3210 0.3902 0.8921 0.5721 0.3670 0.3303 0.0153
Spider [80] ICML 2024 ConvNeXt-L [101] 0.9900 0.4412 0.4740 0.4412 0.6021 0.9314 0.6023 0.4921 0.2344 0.0129
ZoomNeXt [102] TPAMI 2024 ResNet-50 [95] 0.9888 0.5012 0.5079 0.4936 0.6785 0.9508 0.6289 0.5182 0.2161 0.0117
MDCNet [21] CVPR 2024 ResNet-50 [95] 0.9908 0.5652 0.5420 0.5372 0.7084 0.9538 0.6709 0.5690 0.1761 0.0094
MDCNeXt ResNet-50 [95] 0.9913 0.6252 0.6245 0.6611 0.7726 0.9758 0.7459 0.6522 0.1461 0.0086
Tough SAM 2 [78] ICLR 2025 Hiera-B+ [100] 0.9840 0.1312 0.1102 0.0981 0.4980 0.7022 0.5320 0.0921 0.4467 0.0186
SegGPT [76] ICCV 2023 ViT-L [19] 0.9819 0.0914 0.0824 0.0776 0.5092 0.6022 0.5158 0.0790 0.4579 0.0180
DeepLabV3+ [54] ECCV 2018 ResNet-101 [95] 0.9878 0.2712 0.1997 0.1939 0.2892 0.7920 0.3702 0.3083 0.3999 0.0178
SegFormer [67] NeurIPS 2021 MiT-B5 [67] 0.9890 0.2521 0.2001 0.2032 0.2782 0.7808 0.3840 0.3012 0.3911 0.0173
Spider [80] ICML 2024 ConvNeXt-L [101] 0.9896 0.3831 0.3221 0.3353 0.4483 0.9312 0.5201 0.4032 0.2411 0.0146
ZoomNeXt [102] TPAMI 2024 ResNet-50 [95] 0.9888 0.4378 0.3792 0.3511 0.5189 0.9433 0.5811 0.4732 0.2235 0.0129
MDCNet [21] CVPR 2024 ResNet-50 [95] 0.9890 0.4735 0.4012 0.3832 0.5733 0.9511 0.6021 0.5034 0.2190 0.0121
MDCNeXt ResNet-50 [95] 0.9903 0.5641 0.5633 0.5793 0.7353 0.9745 0.7153 0.5872 0.1831 0.0097
Average SAM 2 [78] ICLR 2025 Hiera-B+ [100] 0.9829 0.1884 0.1663 0.1520 0.5245 0.7136 0.5455 0.1421 0.4231 0.0184
SegGPT [76] ICCV 2023 ViT-L [19] 0.9818 0.1472 0.1332 0.1300 0.5284 0.6697 0.5331 0.1316 0.4259 0.0181
DeepLabV3+ [54] ECCV 2018 ResNet-101 [95] 0.9870 0.3717 0.3289 0.3145 0.4297 0.8708 0.5012 0.3707 0.3296 0.0155
SegFormer [67] NeurIPS 2021 MiT-B5 [67] 0.9878 0.3622 0.3210 0.3020 0.4250 0.8584 0.5161 0.3672 0.3384 0.0155
Spider [80] ICML 2024 ConvNeXt-L [101] 0.9896 0.4734 0.4543 0.4517 0.5968 0.9342 0.5950 0.4970 0.2129 0.0127
ZoomNeXt [102] TPAMI 2024 ResNet-50 [95] 0.9888 0.5134 0.4899 0.4810 0.6513 0.9511 0.6290 0.5328 0.1965 0.0115
MDCNet [21] CVPR 2024 ResNet-50 [95] 0.9902 0.5619 0.5256 0.5291 0.6840 0.9567 0.6662 0.5811 0.1738 0.0100
MDCNeXt ResNet-50 [95] 0.9912 0.6232 0.6225 0.6521 0.7699 0.9761 0.7458 0.6489 0.1494 0.0087
V-C Comparison with Image Segmentation Methods
@@
Quantitative Comparison. To comprehensively evaluate MDCNeXt’s segmentation performance, we compare it with generalist models (SAM 2 [78], SegGPT [76]), general segmentation frameworks (DeepLabV3+ [54], SegFormer [67]), and context-dependent concept segmentation models (Spider [80], ZoomNeXt [102]). As shown in Table IV, ten widely adopted segmentation metrics are used for evaluation. From the average scores, SAM 2 and SegGPT perform reasonably on $PA$ and $S_{m}$ due to the dominance of background regions and structure-aware properties, but underperform on other metrics. This suggests that these in-context learning-based models struggle to propagate prompt features into the main segmentation branch for fine-grained guidance, limiting their adaptability to tasks like PBD even with large-scale pretraining. In contrast, Spider and ZoomNeXt benefit from foreground-background comparison and rich-scale fusion, outperforming general segmentation methods. Notably, MDCNeXt achieves the best results across all metrics. On the tough split, it surpasses MDCNet with improvements of 51.2% in $F_{\beta}^{w}$, 18.8% in $mIoU$, and 19.8% in $\mathcal{M}$, showing the effectiveness of our prompt-filtered and density-aware reordering state space modules. Tab. V further presents the efficiency comparison. MDCNeXt still shows clear advantages, achieving a good balance between parameter size and computation.
@@
TABLE V: Efficiency comparison of different segmentation methods in terms of parameter size and computational cost. The best two results are highlighted in red and green, respectively. The worst results are marked with underline.
@@
 [78]  [76]  [54]  [67]  [80]  [102]  [21] (Ours)
@@
TABLE VI: Ablation experiments of each component in the MDCNeXt.
@@
TABLE VII: Ablation experiments of state space modeling, prompt filter and reordering operation in PFSSM and DRSSM. M1: w/o SSM in PSSM. M2: w/o prompt filter in PSSM. M3: w/o reordering operation in DRSSM. M4: replace DRSSM with multiple convolution layers.
@@
TABLE VIII: Quantitative comparison of different label generation strategies and settings for the point segmentation branch. Const-1/3/5 refers to masks with radius of 1/3/5 pixels. The results are averaged across the three test set splits.
@@
TABLE IX: Stability evaluation of MDCNeXt with randomly selected prompt inputs.
@@
TABLE X: Quantitative comparison between joint and separate training for anode and cathode plates.
@@
V-D Ablation Study
Each Component. We analyze the contribution of each component in Tab. VI. Our baseline is a point segmentation branch built on FPN [10] with a ResNet-50 backbone. First, we can see that baseline has been able to outperform most competitors except HS-FPN, validating the suitability of point-based segmentation for the PBD task. Next, we add the counting predictor (CP) and line predictor (LP) to further refine the point segmentation at both high-level and low-level features. The former improves the accuracy of plate counting, while the latter helps adjust the positioning errors. Thus, multi-dimensional features complement each other and improve overall performance. Then, we incorporate the prompt stream to guide the current image stream with pure plate features through simple convolution fusion. This addition provides consistent improvements across all splits under different metrics. Finally, we introduce the state-space modeling into the architecture. PFSSM is integrated into the multi-level feature fusion between current and prompt streams. From the average results, we observe that “+ PFSSM” shows consistent improvements compared to “+ Prompt Stream” with the gain of 14% and 57% in terms of PN-ACC and OH-MAE. DRSSM is added at the tail of the framework to further refine the point maps. On the tough split, PN-ACC increases from 0.5299 to 0.6263, and AL/CL-MAE is reduced by more than 50%.
@@ V-D Ablation Study
Effectiveness of State Space Modeling, Prompt Filter and Reordering Operation. As shown in Tab. VII, we evaluate four variants of MDCNeXt. First, removing the state space modeling in PFSSM (M1) leads to significant performance degradation. Compared to the full MDCNeXt, the average AN-MAE and CN-MAE worsen from 0.4645 and 0.3005 to 0.8847 and 0.5657, while AN-ACC and CN-ACC drop from 0.8705 and 0.8375 to 0.7874 and 0.7144, respectively. These results confirm the critical role of SSM in modeling contextual relationships between the current and prompt streams. Next, removing the prompt filter in PFSSM (M2) leads to a notable performance decline. The average PN-ACC drops from 0.7705 to 0.7062, while OH-MAE deteriorates from 1.0667 to 1.2912. These results indicate that the prompt filter is essential for refining the prompt signal and suppressing irrelevant interference. Then, we verify the effectiveness of the reordering operation by removing it in DRSSM (M3) and replacing DRSSM with multiple convolutional layers (M4). We can see that both modifications lead to consistent performance drops. Thus, simply stacking more convolutional layers cannot match the performance of the proposed reordering-based refinement. Finally, the gap between M3 and M4 suggests that directly applying SSM without the reordering mechanism in the final refinement stage loses its density-aware advantage, even underperforming vanilla convolution.
@@
Label Generation Strategies. For point segmentation, we conduct a series of experiments about label generation strategies as discussed in Sec. IV-E. As shown in Tab. VIII, the model trained under the Ada-0.3 performs best in terms of number accuracy and localization error among all the six settings. The distance-adaptive strategy consistently outperforms the constant strategy.
@@
Selection of Prompts. In Tab. IX, we evaluate the performance of the model using five different groups of P attribute images as the prompt input. It can be seen that MDCNeXt shows strong stability and generalization without dependence on a specific prompt. Thus, the strategy of random selection during training indeed makes MDCNeXt robust against different prompts when testing.
@@
Joint Training vs. Separate Training. Joint training refers to using a single unified model to simultaneously generate predictions related to both anode and cathode plates in a single training process, whereas separate training involves two independent models. Due to the use of a single set of parameters and a joint training process, a mutual learning mechanism can be established between anode and cathode, which helps enhance the model’s discriminative ability and prevents confusion between the two plate types caused by the short overhang. As shown in Tab. X, the jointly trained model achieves superior performance while using only half the parameters compared to the separate training setting.
VI Future Works and Outlook
@@
III) Fine-grained Industrial Image Generation. High-quality and diverse battery data is essential for building robust foundation models in PBD. On the one hand, collecting real X-ray samples from different manufacturers is a long-term effort. On the other hand, we believe that generating targeted synthetic data based on specific interference attributes serves as an important complement, offering significant performance gains for PBD before sufficient real data is available. However, most general-purpose generative models [103, 104] are designed for natural scenes and focus on global content realism, lacking control over intricate local structures, such as individual defects or subtle texture changes. In the future, we plan to develop a controllable fine-grained generation framework tailored to industrial scenarios. Based on end-to-end training with industrial data, the focus will be on achieving pixel-level generation conditioned on interference types and spatial layouts, rather than merely ensuring local and global semantic consistency.
@@
V) High-speed and High-precision Industrial CT Reconstruction. As shown in Fig. 16, we visualize some industrial CT data, including slice views for overhang inspection (a, c) and maximum intensity projection (MIP) images along different axes (b: Z-axis, d: X-axis) for metal inclusions or microcracks. Compared with 2D X-ray imaging, volumetric CT provides more comprehensive structural insights, making it valuable for full-scope battery inspection. We plan to extend the PBD dataset into a 3D version by incorporating CT scans, which will significantly enhance the detection of internal quality issues. However, achieving real-time, high-resolution (>>3K) CT reconstruction under industrial constraints remains highly challenging. Recent advances like 3D gaussian splatting (GS) [105] have shown promise in surface-level reconstruction using multi-view inputs [106, 107]. Future research may explore hybrid neural-physical frameworks that incorporate CT physical projection models and 3D GS representations to enable end-to-end CT reconstruction from arbitrary views. In this process, sparse-view sampling, coarse-to-fine refinement pipelines, and the deployment of customized CUDA-based implementations can help accelerate the reconstruction speed.
VII Conclusion
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62431004 and 62276046.

### v2

@@ 5.2 Comparison with Corner Detection, Crowd Counting, General and Tiny Object Detection Solutions
5.3 Comparison with Image Segmentation Methods
@@
Quantitative Comparison. To comprehensively evaluate MDCNeXt’s segmentation performance, we compare it with generalist models (SAM 2 SAM2 , SegGPT SegGPT ), general segmentation frameworks (DeepLabV3+ DeepLabV3+ , SegFormer SegFormer ), and context-dependent concept segmentation models (Spider CDCU-Spider , ZoomNeXt ZoomNeXt ). As shown in Table 4, ten widely adopted segmentation metrics are used for evaluation. From the average scores, SAM 2 and SegGPT perform reasonably on $PA$ and $S_{m}$ due to the dominance of background regions and structure-aware properties, but underperform on other metrics. This suggests that these in-context learning-based models struggle to propagate prompt features into the main segmentation branch for fine-grained guidance, limiting their adaptability to tasks like PBD even with large-scale pretraining. In contrast, Spider and ZoomNeXt benefit from foreground-background comparison and rich-scale fusion, outperforming general segmentation methods. Notably, MDCNeXt achieves the best results across all metrics. On the tough split, it surpasses MDCNet with improvements of 51.2% in $F_{\beta}^{w}$, 18.8% in $mIoU$, and 19.8% in $\mathcal{M}$, showing the effectiveness of our prompt-filtered and density-aware reordering state space modules. Tab. 5 further presents the efficiency comparison. MDCNeXt still shows clear advantages, achieving a good balance between parameter size and computation.
@@
Table 5: Efficiency comparison of different segmentation methods in terms of parameter size and computational cost. The best two results are highlighted in red and green, respectively. The worst results are marked with underline.
@@
 SAM2  SegGPT  DeepLabV3+  SegFormer  CDCU-Spider  ZoomNeXt  MDCNet (Ours)
@@
Table 6: Ablation experiments of each component in the MDCNeXt.
@@
Table 7: Ablation experiments of state space modeling, prompt filter and reordering operation in PFSSM and DRSSM. M1: w/o SSM in PSSM. M2: w/o prompt filter in PSSM. M3: w/o reordering operation in DRSSM. M4: replace DRSSM with multiple convolution layers.
@@
5.4 Ablation Study
Each Component. We analyze the contribution of each component in Tab. 6. Our baseline is a point segmentation branch built on FPN FPN with a ResNet-50 backbone. First, we can see that the baseline has been able to outperform most competitors except HS-FPN, validating the suitability of point-based segmentation for the PBD task. Next, we add the counting predictor (CP) and line predictor (LP) to further refine the point segmentation at both high-level and low-level features. The former improves the accuracy of plate counting, while the latter helps adjust the positioning errors. Thus, multi-dimensional features complement each other and improve overall performance. Then, we incorporate the prompt stream to guide the current image stream with pure plate features through simple convolution fusion. This addition provides consistent improvements across all splits under different metrics. Finally, we introduce the state-space modeling into the architecture. PFSSM is integrated into the multi-level feature fusion between current and prompt streams. From the average results, we observe that “+ PFSSM” shows consistent improvements compared to “+ Prompt Stream” with the gain of 14% and 57% in terms of PN-ACC and OH-MAE. DRSSM is added at the tail of the framework to further refine the point maps. On the tough split, PN-ACC increases from 0.5299 to 0.6263, and AL/CL-MAE is reduced by more than 50%.
Table 8: Quantitative comparison of different label generation strategies and settings for the point segmentation branch. Const-1/3/5 refers to masks with radius of 1/3/5 pixels. The results are averaged across the three test set splits.
@@ 5.4 Ablation Study
Table 9: Stability evaluation of MDCNeXt with randomly selected prompt inputs.
@@
Table 10: Quantitative comparison between joint and separate training for anode and cathode plates.
@@
Effectiveness of State Space Modeling, Prompt Filter and Reordering Operation. As shown in Tab. 7, we evaluate four variants of MDCNeXt. First, removing the state space modeling in PFSSM (M1) leads to significant performance degradation. Compared to the full MDCNeXt, the average AN-MAE and CN-MAE worsen from 0.4645 and 0.3005 to 0.8847 and 0.5657, while AN-ACC and CN-ACC drop from 0.8705 and 0.8375 to 0.7874 and 0.7144, respectively. These results confirm the critical role of SSM in modeling contextual relationships between the current and prompt streams. Next, removing the prompt filter in PFSSM (M2) leads to a notable performance decline. The average PN-ACC drops from 0.7705 to 0.7062, while OH-MAE deteriorates from 1.0667 to 1.2912. These results indicate that the prompt filter is essential for refining the prompt signal and suppressing irrelevant interference. Then, we verify the effectiveness of the reordering operation by removing it in DRSSM (M3) and replacing DRSSM with multiple convolutional layers (M4). We can see that both modifications lead to consistent performance drops. Thus, simply stacking more convolutional layers cannot match the performance of the proposed reordering-based refinement. Finally, the gap between M3 and M4 suggests that directly applying SSM without the reordering mechanism in the final refinement stage loses its density-aware advantage, even underperforming vanilla convolution.
@@
Label Generation Strategies. For point segmentation, we conduct a series of experiments about label generation strategies as discussed in Sec. 4.5. As shown in Tab. 8, the model trained under the Ada-0.3 performs best in terms of number accuracy and localization error among all the six settings. The distance-adaptive strategy consistently outperforms the constant strategy.
@@
Selection of Prompts. In Tab. 9, we evaluate the performance of the model using five different groups of P attribute images as the prompt input. It can be seen that MDCNeXt shows strong stability and generalization without dependence on a specific prompt. Thus, the strategy of random selection during training indeed makes MDCNeXt robust against different prompts when testing.
@@
Joint Training vs. Separate Training. Joint training refers to using a single unified model to simultaneously generate predictions related to both anode and cathode plates in a single training process, whereas separate training involves two independent models. Due to the use of a single set of parameters and a joint training process, a mutual learning mechanism can be established between anode and cathode, which helps enhance the model’s discriminative ability and prevents confusion between the two plate types caused by the short overhang. As shown in Tab. 10, the jointly trained model achieves superior performance while using only half the parameters compared to the separate training setting.
6 Future Works and Outlook
@@
III) Fine-grained Industrial Image Generation. High-quality and diverse battery data is essential for building robust foundation models in PBD. On the one hand, collecting real X-ray samples from different manufacturers is a long-term effort. On the other hand, we believe that generating targeted synthetic data based on specific interference attributes serves as an important complement, offering significant performance gains for PBD before sufficient real data is available. However, most general-purpose generative models Stable_diffusion ; ControlNet are designed for natural scenes and focus on global content realism, lacking control over intricate local structures, such as individual defects or subtle texture changes. In the future, we plan to develop a controllable fine-grained generation framework tailored to industrial scenarios. Based on end-to-end training with industrial data, the focus will be on achieving pixel-level generation conditioned on interference types and spatial layouts, rather than merely ensuring local and global semantic consistency.
@@
V) High-speed and High-precision Industrial CT Reconstruction. As shown in Fig. 17, we visualize some industrial CT data, including slice views for overhang inspection (a, c) and maximum intensity projection (MIP) images along different axes (b: Z-axis, d: X-axis) for metal inclusions or microcracks. Compared with 2D X-ray imaging, volumetric CT provides more comprehensive structural insights, making it valuable for full-scope battery inspection. We plan to extend the PBD dataset into a 3D version by incorporating CT scans, which will significantly enhance the detection of internal quality issues. However, achieving real-time, high-resolution (>>3K) CT reconstruction under industrial constraints remains highly challenging. Recent advances like 3D gaussian splatting (GS) 3DGS have shown promise in surface-level reconstruction using multi-view inputs Splatflow ; VGGT . Future research may explore hybrid neural-physical frameworks that incorporate CT physical projection models and 3D GS representations to enable end-to-end CT reconstruction from arbitrary views. In this process, sparse-view sampling, coarse-to-fine refinement pipelines, and the deployment of customized CUDA-based implementations can help accelerate the reconstruction speed.
7 Conclusion
@@
Data availability statement: All datasets used and studied in this paper are publicly available. The source code and datasets are publicly available at PBD5K.